OpenStudy (anonymous):

ok guys tell me if i got this one right solve 2x^2=10 find the x intercepts of f(x)= 2x^2-10

OpenStudy (anonymous):

solutions are \[(\sqrt{5},-\sqrt{5})\]

OpenStudy (anonymous):

intercepts are (5,0),(-5,0)

OpenStudy (chaise):

The first answers are correct. Your intercepts are wrong. When you solve for x you get -sqrt(5) and sqrt(5). y =2x^2-10 When finding the x intercept, y is 0. (which is what you got in the first question) When finding the y intercept x = 0 y = 2x^2 -10 y = 2x0^2 -10 y = -10 Your intercepts are \[(\sqrt{5}, 0) and (\sqrt{-5},0)\]

OpenStudy (chaise):

That's your x intercepts, and your y intercept is -10